sophiaaa:

A table contains the ordered pairs (3, 42.25) and (5, 50.75). If the relationship in the table is linear, explain how to find the initial value.

swinn:

The linear relationship (the linear equation): y = m x + b (the slope-intercept form) m = (50.75 - 42.25) / (5 - 3) = 8.50 / 2 = 4.25 (the slope) 42.25 = 4.25 * 3 + b b = 42.25 - 12.75 b = 29.50 y = 4.25 * x + 29.50 initial value= Hope this helps:)

sophiaaa:

Thank you so much and it was correct :>>>

swinn:

np